OpenStudy (anonymous):

The length of a rectangle is 2 more than the width. The area is 15 square inches. Find the dimensions

OpenStudy (anonymous):

L = W + 2 L*W = 15 Solve

OpenStudy (anonymous):

You got it?

OpenStudy (anonymous):

then if you really don't know how to solve it the second step is substitute the value of L to the second equation then solve for W.

OpenStudy (anonymous):

i did: w+2+w=15. is that wrong? i keep getting bad answers

OpenStudy (anonymous):

then the nxt is 2w= 15-2

OpenStudy (anonymous):

W(W+2)=15 W^2 + 2W = 15 W^2 + 2W - 15 = 0 (W+5)(W-3) = 0 W = -5 or 3. Since it can't be negative: W = 3 L = 5